# Best of both worlds: Fusing hyperspectral data from two generations of spectro-imagers for X-ray astrophysics

### v1

@@ Abstract
Context. With the recent launch of the X-Ray Imaging and Spectroscopy Mission (XRISM) and the advent of microcalorimeter detectors, X-ray astrophysics is entering in a new era of spatially resolved high resolution spectroscopy. But while this new generation of X-ray telescopes have much finer spectral resolutions than their predecessors (e.g. XMM-Newton, Chandra), they also have coarser spatial resolutions, leading to problematic cross-pixel contamination. This issue is currently a critical limitation for the study of extended sources such as galaxy clusters of supernova remnants.
Aims. To increase the scientific output of XRISM’s hyperspectral data, we propose to fuse it with XMM-Newton data, and seek to obtain a cube with the best spatial and spectral resolution of both generations. This is the aim of hyperspectral fusion. In this article, we explore the potential and limitations of hyperspectral fusion for X-ray astrophysics.
Methods. We have implemented an algorithm that jointly deconvolves the spatial response of XRISM and the spectral response of XMM-Newton. To do so, we construct a forward model adapted for instrumental systematic degradations and Poisson noise, then tackle hyperspectral fusion as a regularized inverse problem. We test three methods of regularization: spectral low rank approximation with a spatial Sobolev regularization; spectral low rank approximation with a 2D wavelt sparsity constraint ; and a 2D-1D wavelet sparsity constraint.
Results. We test our method on toy models constructed from hydrodynamic simulations of supernova remnants. We find that our method reconstructs the ground truth well even when the toy model is complex. For the regularization term, we find that while the low rank approximation worked well as a spectral denoiser in models with less spectral variability, it introduced a bias in models with more spectral variability, in which case the 2D-1D wavelet sparsity regularization worked best. After demonstrating a proof of concept in this article, we aim to apply this method to real X-ray astrophysical data in the near future.
@@ 1 Introduction
With the advent of microcalorimeter detectors, X-ray astrophysics has entered in a new era of spatially resolved high resolution spectroscopy. Indeed, the recent launch of the X-ray observatory XRISM and its microcalorimeter array Resolve [^1] https://xrism.isas.jaxa.jp, and the future Athena telescope and its XIFU [^2] https://x-ifu.irap.omp.eu/ camera, are beginning to grant access to spatially resolved spectroscopy at spectral resolutions improved by an order of magnitude. This will allow astrophysicists to investigate previously indistinguishable spectral features, such as the emission lines of rare elements, line broadening, or velocity measurements. These characteristics are essential in the study of extended sources such as supernova remnants, galaxy clusters, and many of the other exotic sources of the X-ray universe.
However, while the XRISM/Resolve instrument has a much finer spectral resolution than Chandra/ACIS [^3] https://cxc.harvard.edu/ or XMM-Newton/EPIC [^4] https://www.cosmos.esa.int/web/xmm-newton cameras, its spatial resolution is much lesser than that of its predecessors. As we will see, this spatial blur is an important obstacle when analyzing XRISM/Resolve data. Given this trade-off, could one combine the data from these two generations of X-ray observatories and obtain a result with the best spectral and spatial resolutions of each instrument? This is the idea at the heart of hyperspectral fusion. In this paper, we explore the potential and limitations of hyperspectral fusion for X-ray astrophysics, and test our method on simulations resembling a supernova remnant’s observations.
@@ 1.1 Astrophysical context: X-ray spectro-imaging and microcalorimeters
X-ray spectro-imagers can collect hyperspectral images. These are three-dimensional data arrays: they have two spatial dimensions, like a classical image, and a third spectral dimension (energy, or wavelength). That is to say, to each spatial pixel is associated an energy spectrum.
@@
In the last decades, Charge-coupled device (CCD) cameras have been the main spectro-imaging instruments onboard X-ray telescopes such as XMM-Newton, Chandra, Swift, Suzaku and eROSITA. When the X-ray photon interacts with the semiconductor material, an electric charge is generated proportional to the energy of the incident photon. The charge is then transferred row by row to a readout register. For example, the XMM-Newton MOS camera can accommodate a large number (600$\times$600) of small pixels (40 $\mu$m, Turner et al., 2001). The energy resolution is of the order 150 eV [^5] https://xmm-tools.cosmos.esa.int/external/xmm_user_support/documentation/uhb/epic_specres.html at 6 keV ($E\,/\Delta E\sim 40$) which is close to the Fano limit, the physical lower bound of energy resolution due to the statistical fluctuations in the number of electron-hole pairs produced in the photon-material interaction.
@@
To overcome this limit, a change of technology is required. Microcalorimeter devices are providing this revolutionary leap forward for X-ray spectro-imaging instruments. The fundamental principle of microcalorimeters is to measure the photon energy via the small increase in temperature caused by the conversion of this energy into heat in the sensor. To be sensitive to small heat increments and reach a spectral resolution in the order of a few eV, the device has to be cooled down to 50 mK with very stable temperature. It is important to note that, unlike gratings onboard XMM-Newton or Chandra, microcalorimeters are non-dispersive spectrometers and can be used to study diffuse structure.
@@
However, with the current available technology, only a limited number of pixels can be accommodated on the focal plane. The XRISM/Resolve instrument is a microcalorimeter array of 6 × 6 sensors (pixel size of 800 $\mu$m, Takahashi et al., 2014) with a constant spectral resolution over the entire bandpass of 5 eV FWHM (XRISM Science Team, 2020). While future X-ray telescopes combining microcalorimeter and high spatial resolution are planned (Athena/XIFU, see Barret et al., 2018, for more details), they are expected to launch in the late 2030’s, and would still not reach the spatial resolution of the X-ray telescope Chandra.
@@
In the meantime, mitigation plans to overcome this spatial resolution limitation should be investigated to try to partially recover some of the spatial information. To that end, we explore in this study the performance of hyperspectral fusion to combine data sets with high spectral resolution and high spatial resolution, using the example of fusion between XMM-Newton and XRISM/Resolve data. The final goal of such reconstruction is to enhance the scientific return of XRISM/Resolve by retrieving maps of physical parameters (plasma temperature, velocity, abundances, etc) at higher resolution with reduced cross-contamination betweeen pixels (spectrospatial mixing).
@@
Indeed, this mixing is an important problem when analysing hyperspectral data. If an instrument’s spatial resolution is larger than the coherence scale of a physical parameter, the reconstructed information will be an average of the small scale features. To take the example of a velocity map from the ejecta of a supernova remnant, the small scale redshifted/blueshifted ejecta features can mistakenly be measured at small or close to zero redshift values, as was shown in Godinaud et al. (2023)’s detailed analysis of Tycho’s supernova remnant. By jointly deconvolving the data sets from two X-ray instruments, hyperspectral data fusion could solve this problem and recover a higher fidelity scientific information at small scale.
@@ 1.2 Hyperspectral Fusion
Given two hyperspectral (HS) images, the general task of hyperspectral fusion consists in obtaining a cube with the best spatial and spectral resolution of each data set. For instance, Fig. 1 shows HS images of the Perseus galaxy cluster taken respectively by the X-ray telescopes XMM-Newton and Hitomi/SXS [^6] XRISM/Resolve camera is identical to the SXS onboard the short-lived Hitomi telescope which was launched in 2016.. XMM-Newton has better spatial resolution, and Hitomi has better spectral resolution. The goal of fusing the two datasets would be to obtain a HS image with XMM-Newton’s spatial resolution and Hitomi’s spectral resolution.
This problem has been explored in the literature for applications to satellite pictures of the earth, and has recently received some attention in astrophysics with the launch of the James Webb Space Telescope (JWST). However, X-ray astrophysics has its own specific challenges. X-ray data is dominated by Poisson noise, and both instruments produce HS images with many spectral channels, as opposed to one of the data sets being a multispectral (MS) image with a small amount of spectral bands typically obtained with different filters. To use hyperspectral fusion for X-ray astrophysics, it is thus key to find the fusion method most appropriate for this case study.
In this section, we give an overview of the various aspects to take into account when approaching a problem of hyperspectral fusion, in order to find the best way to solve the problem. For a detailed review of the literature, we refer the interested reader to Yokoya et al. (2017).
The first aspect to consider is the forward model of the two instruments. This defines how a source image will be degraded in the process of image formation and recording. It will depend on the systematic characteristics of the instrument, and the noise we expect to measure.
For the systematic instrumental degradation, two dimensions may be taken into account: the spatial degradation, and the spectral degradation. Spatially, the degradation will be due to optical diffraction or other systematic blurs, and this information is encapsulated in the point source function (PSF). Given a point source, the PSF returns its radiance distribution on the recorded image. Similarly, if a source emits a Dirac spectra, the measured spectra will be the instrument’s spectral response function, and this will be affected by, for instance, the chemical and physical properties of the spectro-imager. Depending on the instrument, the PSF and spectral response may depend on the energy, on the spatial location of the pixel on the camera, or both.
@@
Once the instrumental systematic degradation and the stochastic noise have been taken into account, if we take two HS images $X$ and $Y$ of sizes ($M\times N\times l$) and ($m\times n\times L$), the general problem to fuse two HS images will be:
@@
where $\hat{Z}$ is the fusion result, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ are the model operators that take into account the systematic instrumental degradation (PSF, spectral response…) and the rebinning needed to take account different voxel sizes, and $D$ is a well-chosen function that evaluates the difference between the model $\mathcal{M}_{X}(Z)$ and data $X$, and between model $\mathcal{M}_{Y}(Z)$ and data $Y$, given the expected noise distribution. Thus, this problem ensures that $Z$ is a plausible source from which the first instrument would measure $X$, and the second would measure $Y$. In the case of X-ray astrophysics, the forward model is the one described in section 2.1 and Fig. 2.
However, this problem is often ill-posed (in the sense of Hadamard, 1923) and ill-conditioned, meaning that while there might be a unique solution, the noise will be amplified when inverting the model, corrupting the solution, and slight changes in the data would lead to drastic changes in the solution (Bertero M. & C., 2022). One well-studied approach to solve ill-posed inverse problems is to regularize the problem, which restricts the space of acceptable solutions. This can be approached by considering the spatial features of $Z$, its spectral features, or both. To regularize in the spectral dimension, one classic method is to project the spectra of $Z$ onto a lower-dimensional subspace, such as is done with source separation (e.g. Yokoya et al., 2012; Prévost et al., 2022; Pineau et al., 2023) or low-rank approximation (e.g. Simões et al., 2015; Xu et al., 2022; Guilloteau et al., 2020). We can also add a regularization term $\varphi(Z)$ to the function we seek to minimize, which leads us to:
@@
This $\varphi(Z)$ will promote certain spatial or spectral structures in the solution. Its choice must thus be appropriate for our data. For instance, in remote sensing (i.e. the study of satellite pictures of the earth), it may be useful to promote spatial piece-wise smoothness using a total variation term such as done by Simões et al. (2015), but this would be ill-advised for astrophysical data, where such features are rarely observed. For our case of X-ray astrophysics, we consider three potential regularization approach, which are described in section 2.3.
Finally, to solve the problem in equation 2, we need to pick an appropriate optimization scheme. In some cases, there might be an analytical solution (such as Pineau et al. (2023)’s work for JWST, when the spectral content is assumed to be a linear combination of known spectra), but most often an iterative algorithm will have to be constructed. The choice of optimizer will depend on the data size, the specific challenges presented by the forward model, and the chosen regularization terms. It may also be the case that constraints placed on the optimizer (such as, wishing that the algorithm reaches convergence as quickly as possible) will lead to constraints on the choice of regularization and cost function.
Figure 2: A simplified diagram of the forward model of two X-ray telescopes (here, XMM-Newton/EPIC and XRISM/Resolve). The problem of fusion consists in inverting this model, using the observed data X and Y to retrieve the deconvoluted cube Z. The mathematical details behind the operators and kernels are detailed in section 2. Not shown is the exposure time of each data sets.
1.3 Contributions
In astrophysics, work has been done to tackle MS/HS fusion for the JWST (see Guilloteau et al., 2022; Pineau et al., 2023), to fuse images taken by two of its instruments (NIRCam and NIRSpec), but to our knowledge, no algorithm has been implemented for HS/HS fusion with Poisson statistics, considering two instruments with their PSF and spectral responses, as is needed in X-ray astrophysics.
In this work, we investigate a hyperspectral fusion method for X-ray astrophysics. Our aim is to evaluate the potential and limitations of fusion for X-ray data taken by couples of telescopes such as XMM-Newton/EPIC with XRISM/Resolve, or Chandra with Athena-XIFU. Section 2 presents our methodology, going over the forward model we have developed for X-ray astrophysics, and the three potential regularization schemes we consider: first, inspired by Guilloteau et al. (2020), and given the widespread usage of low-rank approximations in the literature, we implement a low-rank approximation with Sobolev spatial regularization. Since Sobolev regularization does not take into account multi-scale features, we then consider a low-rank approximation with a 2D wavelet sparsity regularization. Finally, we investigate whether low-rankedness is indeed appropriate, and also implement a heuristic approach using a 2D-1D Wavelet sparsity regularization.
We have implemented a modular optimization algorithm which can use either of these three potential regularizations. Its code is openly available on Github. [^7] https://github.com/JMLascar/HIFReD_fusion In section 3, we test our method on toy models of varying complexity, using synthetic X-ray data cubes constructed from hydrodynamic simulation of supernova remnants. We discuss the achievements, limitations, and perspective for future work in section 4.

### v2

@@ Abstract
Context. With the recent launch of the X-Ray Imaging and Spectroscopy Mission (XRISM) and the advent of microcalorimeter detectors, X-ray astrophysics is entering in a new era of spatially resolved high-resolution spectroscopy. But while this new generation of X-ray telescopes offers much finer spectral resolution than the previous one (e.g. XMM-Newton, Chandra), these instruments also have coarser spatial resolutions, leading to problematic cross-pixel contamination. This issue is currently a critical limitation for the study of extended sources, such as galaxy clusters or supernova remnants (SNRs).
Aims. To increase the scientific output of XRISM’s hyperspectral data, we propose that it be fused with XMM-Newton data, and seek to obtain a cube with the best spatial and spectral resolution of both generations. This is the aim of hyperspectral fusion. In this article, we explore the potential and limitations of hyperspectral fusion for X-ray astrophysics.
Methods. We implemented an algorithm that jointly deconvolves the spatial response of XRISM and the spectral response of XMM-Newton. To do so, we construct a forward model adapted for instrumental systematic degradations and Poisson noise, and then tackle hyperspectral fusion as a regularized inverse problem. We test three methods of regularization: spectral low-rank approximation with a spatial Sobolev regularization; spectral low-rank approximation with a 2D wavelet sparsity constraint; and a 2D-1D wavelet sparsity constraint.
Results. We test our method on toy models constructed from hydrodynamic simulations of SNRs. We find that our method reconstructs the ground truth well even when the toy model is complex. For the regularization term, we find that while the low-rank approximation works well as a spectral denoiser in models with less spectral variability, it introduces a bias in models with more spectral variability, in which case the 2D-1D wavelet sparsity regularization works best. Following the present proof of concept, we aim to apply this method to real X-ray astrophysical data in the near future.
@@ 1 Introduction
With the advent of microcalorimeter detectors, X-ray astrophysics has entered a new era of spatially resolved high-resolution spectroscopy. Indeed, the recent launch of the X-Ray Imaging and Spectroscopy Mission (XRISM) and its microcalorimeter array ”Resolve” [^1] https://xrism.isas.jaxa.jp, and the future Athena telescope and its X-ray Integral Field Unit (X-IFU)  [^2] https://x-ifu.irap.omp.eu/ camera, has granted access to spatially resolved spectroscopy at spectral resolutions improved by an order of magnitude. This is set to allow astrophysicists to investigate previously indistinguishable spectral features, such as the emission lines of rare elements, line broadening, or velocity measurements. These characteristics are essential in the study of extended sources such as supernova remnants (SNRs), galaxy clusters, and many of the other exotic sources of the X-ray Universe.
However, while the XRISM/Resolve instrument has a much finer spectral resolution than the Chandra X-ray Observatory’s Advanced CCD Imaging Spectrometer (Chandra/ACIS) [^3] https://cxc.harvard.edu/ or the X-ray Multi-Mirror-Newton’s European Photon Imaging Camera (XMM-Newton/EPIC) [^4] https://www.cosmos.esa.int/web/xmm-newton cameras, its spatial resolution is much lower than that of its predecessors. As we demonstrate here, this spatial blur is an important obstacle when analyzing XRISM/Resolve data. Given this trade-off, we set out to establish whether or not it is possible to combine the data from these two generations of X-ray observatories and obtain a result with the best spectral and spatial resolution of each instrument. This is the idea at the heart of hyperspectral fusion.
@@
In this paper, we explore the potential and limitations of hyperspectral fusion for X-ray data taken by pairs of telescopes such as XMM-Newton/EPIC with XRISM/Resolve, or Chandra with Athena-X-IFU.
We implemented a modular optimization algorithm that can use one of three potential regularizations: first, inspired by Guilloteau et al. (2020), and given the widespread usage of low-rank approximations in the literature, we implement a low-rank approximation with Sobolev spatial regularization. Since Sobolev regularization does not take into account multi-scale features, we then considered a low-rank approximation with a 2D wavelet sparsity regularization. Finally, to investigate whether low-rankedness is indeed appropriate, we implemented a heuristic approach using a 2D-1D Wavelet sparsity regularization. This hyperspectral fusion code is openly available on Github. [^5] https://github.com/JMLascar/HIFReD_fusion
In Section 2, we present the astrophysical and methodological context behind our work. In Section 3 we present our methodology, and describe the forward model we have developed for X-ray astrophysics, and the proposed regularization schemes. In section 4, we test our method on toy models of varying complexity using synthetic X-ray data cubes constructed from hydrodynamic simulation of SNRs. We discuss the achievements, limitations, and perspective for future work in section 5.
2 Context
@@ 2.1 Astrophysical context: X-ray spectro-imaging and microcalorimeters
X-ray spectro-imagers can collect hyperspectral images. These are 3D data arrays: they have two spatial dimensions, like a classical image, and a third spectral dimension (energy or wavelength). In other words, an energy spectrum is associated to each spatial pixel.
@@
In recent decades, charge-coupled device (CCD) cameras have been the main spectro-imaging instruments on board X-ray telescopes such as XMM-Newton, Chandra, Swift, Suzaku, and eROSITA. When the X-ray photon interacts with the semiconductor material, an electric charge is generated proportional to the energy of the incident photon. The charge is then transferred row by row to a readout register. For example, the XMM-Newton MOS camera can accommodate a large number (600$\times$600) of small pixels (40 $\mu$m, Turner et al., 2001). The energy resolution is of the order 150 eV [^6] https://xmm-tools.cosmos.esa.int/external/xmm_user_support/documentation/uhb/epic_specres.html at 6 keV ($E\,/\Delta E\sim 40$), which is close to the Fano limit, the physical lower bound of energy resolution due to the statistical fluctuations in the number of electron-hole pairs produced in the photon-material interaction.
@@
To overcome this limit, a change of technology is required. Microcalorimeter devices are providing this revolutionary leap forward for X-ray spectro-imaging instruments. The fundamental principle of microcalorimeters is to measure the photon energy via the small increase in temperature caused by the conversion of this energy into heat in the sensor. To be sensitive to small heat increments and reach a spectral resolution on the order of a few electron volts, the device has to be cooled down to 50 mK and remain at very stable temperature. It is important to note that, unlike gratings on board XMM-Newton or Chandra, microcalorimeters are non-dispersive spectrometers and can be used to study diffuse structure.
@@
However, with the current available technology, only a limited number of pixels can be accommodated on the focal plane. The XRISM/Resolve instrument is a microcalorimeter array of 6 × 6 sensors (pixel size of 800 $\mu$m, Takahashi et al., 2014) with a constant spectral resolution over the entire bandpass of 5 eV FWHM (XRISM Science Team, 2020). While future X-ray telescopes combining microcalorimeter and high spatial resolution are planned (Athena/XIFU, see Barret et al., 2018, for more details), they are expected to launch in the late 2030s, and will still not reach the spatial resolution of the X-ray telescope Chandra.
@@
In the meantime, mitigation plans to overcome this spatial-resolution limitation should be investigated to try to partially recover some of the spatial information. To that end, we explore in this study the performance of hyperspectral fusion to combine data sets with high spectral resolution and high spatial resolution using the example of fusion between XMM-Newton and XRISM/Resolve data. The final goal of this reconstruction is to enhance the scientific return of XRISM/Resolve by retrieving maps of physical parameters (plasma temperature, velocity, abundances, etc.) at higher resolution with reduced cross-contamination between pixels (spectrospatial mixing).
@@
Indeed, this mixing is an important problem when analyzing hyperspectral data. If an instrument’s spatial resolution is larger than the coherence scale of a physical parameter, the reconstructed information will be an average of the small-scale features. To take the example of a velocity map from the ejecta of a SNR, the small-scale redshifted and blueshifted ejecta features can mistakenly be measured at small or close-to-zero redshift values, as was demonstrated by the detailed analysis of Tycho’s SNR in Godinaud et al. (2023). By jointly deconvolving the data sets from two X-ray instruments, hyperspectral data fusion could solve this problem and recover higher-fidelity scientific information at small scale.
@@ 2.2 Hyperspectral fusion
Given two hyperspectral (HS) images, the general task of hyperspectral fusion consists in obtaining a cube with the best spatial and spectral resolution of each data set. For instance, Fig. 1 shows HS images of the Perseus galaxy cluster taken by the X-ray telescopes XMM-Newton and Hitomi/SXS [^7] XRISM/Resolve camera is identical to the SXS on board the short-lived Hitomi telescope launched in 2016.. XMM-Newton has better spatial resolution, and Hitomi has better spectral resolution. The goal of fusing the two datasets would be to obtain a HS image with XMM-Newton’s spatial resolution and Hitomi’s spectral resolution.
This problem has been explored in the literature for applications to satellite pictures of the earth, and has recently received some attention in astrophysics with the launch of the James Webb Space Telescope (JWST). However, X-ray astrophysics has its own specific challenges. X-ray data are dominated by Poisson noise, and both instruments produce HS images with many spectral channels, as opposed to one of the data sets being a multi-spectral image with a small number of spectral bands typically obtained with different filters. To use hyperspectral fusion for X-ray astrophysics, it is thus important to find the fusion method most appropriate for this case study.
In this section, we provide an overview of the various aspects to take into account when approaching a problem of hyperspectral fusion in order to find the best way to solve the problem. For a detailed review of the literature, we refer the interested reader to Yokoya et al. (2017).
The first aspect to consider is the forward model of the two instruments. This defines how a source image will be degraded in the process of image formation and recording, which depends on the systematic characteristics of the instrument, and the noise we expect to measure.
When considering systematic instrumental degradation, two dimensions may be taken into account: the spatial degradation, and the spectral degradation. Spatially, the degradation will be due to optical diffraction or other systematic blurs, and this information is encapsulated in the point spread function (PSF). Given a point source, the PSF returns its radiance distribution on the recorded image. Similarly, if a source emits a Dirac spectrum, the measured spectrum will be the instrument’s spectral response function, and this will be affected by, for instance, the chemical and physical properties of the spectro-imager. Depending on the instrument, the PSF and spectral response may depend on the energy, on the spatial location of the pixel on the camera, or both.
@@
Once the instrumental systematic degradation and the stochastic noise have been taken into account, if we take two HS images $X$ and $Y$ of sizes ($M\times N\times l$) and ($m\times n\times L$), the general problem of fusing two HS images can be written as:
@@
where $\hat{Z}$ is the fusion result, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ are the model operators that take into account the systematic instrumental degradation (PSF, spectral response, etc.) and the rebinning needed to take account different voxel sizes, and $D$ is a well-chosen function that evaluates the difference between the model $\mathcal{M}_{X}(Z)$ and data $X$, and between model $\mathcal{M}_{Y}(Z)$ and data $Y$, given the expected noise distribution. Thus, this problem ensures that $Z$ is a plausible source from which the first instrument can measure $X$ and the second can measure $Y$. In the case of X-ray astrophysics, the forward model is the one described in section 3.1 and Fig. 2.
However, this problem is often ill-posed (in the sense of Hadamard, 1923) and ill-conditioned, meaning that while there might be a unique solution, the noise will be amplified when inverting the model, corrupting the solution, and slight changes in the data would lead to drastic changes in the solution (Bertero M. & C., 2022). One well-studied approach to solving ill-posed inverse problems is to regularize the problem, which restricts the space of acceptable solutions. This can be approached by considering the spatial features of $Z$, its spectral features, or both. To regularize in the spectral dimension, one classic method is to project the spectra of $Z$ onto a lower-dimensional subspace, as is done with source separation (e.g., Yokoya et al., 2012; Prévost et al., 2022; Pineau et al., 2023) or low-rank approximation (e.g., Simões et al., 2015; Xu et al., 2022; Guilloteau et al., 2020). We can also add a regularization term $\varphi(Z)$ to the function we seek to minimize, which leads us to:
@@
This $\varphi(Z)$ will promote certain spatial or spectral structures in the solution. Its choice must thus be appropriate for our data. For instance, in remote sensing (i.e., the study of satellite pictures of the earth), it may be useful to promote spatial piece-wise smoothness using a total variation term, as done by Simões et al. (2015), but this would be ill-advised for astrophysical data, where such features are rarely observed. For our case of X-ray astrophysics, we consider three potential regularization approaches, which are described in section 3.3.
Finally, to solve the problem in equation 2, we need to choose an appropriate optimization scheme. In some cases, there might be an analytical solution (as in the work of Pineau et al. (2023) for JWST, when the spectral content is assumed to be a linear combination of known spectra), but most often an iterative algorithm must be constructed. The choice of optimizer will depend on the data size, the specific challenges presented by the forward model, and the chosen regularization terms. It may also be the case that constraints placed on the optimizer (such as insisting that the algorithm reaches convergence as quickly as possible) will lead to constraints on the choice of regularization and cost function.
Figure 2: Simplified diagram of the forward model of two X-ray telescopes (here, XMM-Newton/EPIC and XRISM/Resolve). The problem of fusion consists in inverting this model and using the observed data X and Y to retrieve the deconvoluted cube Z. The mathematical details behind the operators and kernels are detailed in section 3. Not shown is the exposure time of each data set.
In astrophysics, work has been done to tackle multi-spectral/HS fusion for the JWST (see Guilloteau et al., 2022; Pineau et al., 2023) in order to fuse images taken by two of its instruments (NIRCam and NIRSpec), but to our knowledge, no algorithm has been implemented for HS/HS fusion with Poisson statistics, considering two instruments with their PSF and spectral responses, as is needed in X-ray astrophysics.